# LLM-SAP: Large Language Model Situational Awareness Based Planning

### v3

@@ Abstract
This work pioneers evaluating emergent planning capabilities based on situational awareness in large language models. We contribute (i) novel benchmarks and metrics for standardised assessment; (ii) a unique dataset to spur progress; and (iii) demonstrations that are prompting and multi-agent schemes significantly enhance planning performance in context-sensitive planning tasks. Positioning this within a situated agent and automated planning research, we highlight inherent reliability challenges–efficiently mapping world states to actions without environmental guidance remains open despite simulated domain advances. Although out-of-scope, limitations around validation methodology and data availability indicate exciting directions, including fine-tuning on expanded planning corpora and optimizations for triggering fast latent planning. By conclusively demonstrating current methods’ promise and limitations via rigorous comparison, we catalyze investigating reliable goal-directed reasoning for situated agents. The dataset is available at here.
Index Terms—  Large Language Model, Planning, Situational Awareness, Situated Agent, Multi-Agent Reasoning
@@ 1 Introduction
Developing AI agents that can make flexible decisions is challenging, as they need to handle unpredictability in the real world [1]. Humans determine appropriate interventions through situational awareness, while inadequate situation awareness has been identified as a primary cause of accidents due to human error [2, 3]. As Yadav [4] notes, rigorously examining situational awareness in large language models (LLMs) is crucial to steer their safe and reliable development. Without broader awareness, any seemingly beneficial action risks unintended harm [5, 6]. For example, an autonomous agent may need nuanced judgements to avoid potential harm in situations like a toddler reaching for a pot or playing with a knife.
@@
This research shows LLMs can exhibit human-like planning capacities based on situational awareness. As Fig. 1 illustrates, when prompted for situational awareness based planning spanning perception, comprehension and projection [2] or when provided reasoning feedback, the LLM displays improved deductive reasoning requiring perspective-taking and contemplating potential outcomes. This study distinguishes itself from previous AI reasoning and planning research in significant ways. Much existing work evaluates task-specific planning or makes assumptions about predefined steps under strict conditions [7, 8, 9, 10, 11, 12]. Additionally, various agents reactively generate plans only after explicit instructions or demands [13, 14, 15, 16, 17, 18, 19, 20], rather than proactively. In contrast, our approach centres on assessing and enhancing models’ capacities for proactive situational planning when faced with real, open-world dilemmas. Without rigid constraints or environmental feedback, the models must leverage deductive logic to map conjectured actions while weighing consequences using only a descriptive starting scenario and prompts. By prioritizing latent planning that mirrors human cognition, namely situational awareness based rather than domain-limited planning, our method signifies a break from former suppositions and appraisals.
The remainder of this paper overviews the methodology in Section 2, details experiments in Section 3, analyzes results in Section 4 and concludes with a summary and future directions in Section 5.
@@ 2.1 Task Formulation and Key Challenges
We formulate situational awareness based planning as grounded inference over a dynamic hazard scenario s, with $s\in S$ where S denotes the hazard situation space. The input comprises an unordered set of concepts $x=\{c1,c2,...,ck\}\subseteq C$ encapsulating entities, events, and temporal evolution within s, with C the overall conceptual vocabulary. The output is a step-wise plan $\pi:S\to A$ with actions $a1,a2,...\in A$, where A denotes the action space of possible interventions. Learning the planning policy $\pi:S\to A$ requires overcoming two intrinsic challenges:
  • Failing to achieve high levels of situational awareness - lacking perception, comprehension, and projection of the hazard environment when deducing appropriate state-action mappings
  • Struggling to anticipate potential downstream consequences of planned actions on human safety and property due to inadequate understanding of hazardous situation dynamics
@@ 2.2 Multi-AI Agents Enhance Reasoning and Accuracy
Recent work has shown that employing multiple LLMs in a cooperative framework, whether collaborative or adversarial, can enhance reasoning and factuality accuracy. As Du et al. discuss, the debate between agents allows them to critique each other’s logic, correcting flaws [21]. Similarly, Liang et al. find disagreement motivates broader reasoning as agents try to outdo each other [22]. In these cases, collaboration complements individual strengths [21, 22, 23]. In this work, we use two LLM agents - $\text{LLM}_{gen}$ for plan generation and $\text{LLM}_{eval}$ for critical evaluation. We rely on the closed-loop promotion between these complementary roles to improve latent planning.
@@ 2.3 State-based Planning with Feedback
Current AI systems that rely solely on rigid, context-insensitive rules risk unintended outcomes when deployed in complex, real-world environments [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. To enable more reliable and ethical decision-making, architectures should aspire to model interdependent variables and causal relationships, similar to fluid human reasoning processes.
One approach is to have LLM agents iteratively generate and assess potential solutions before implementation. As an example, we frame the design of a finite state machine (FSM) [24] as a collaborative process between two models. A latent FSM plan can be defined by a tuple $M=(S,T,A)$, comprising a set of states $S$, transitions $T$, and actions $A$. The process begins by representing the plan’s reasoning process as $R$, with a generator agent reasoning ($R_{LLM_{gen}}$) to plan then an evaluator agent reasoning ($R_{LLM_{eval}}$) to evaluate. $R_{LLM_{gen}}$ proposes a candidate FSM plan $\hat{M}$, which $R_{LLM_{eval}}$ then scores and provides feedback $f$ on. $R_{LLM_{gen}}$ incorporates this feedback into the next proposal. This iterative loop continues until the score of $\hat{M}$ is higher than that of $M^{*}$(the benchmark plan), which is finally adopted as the new optimal plan $M^{*}$. This approach, shown in Algorithm 1, allows for tight refinement loops resembling human reasoning. By assessing solutions before real-world implementation, unintended outcomes can potentially be anticipated and mitigated.
Algorithm 1 Situational Awareness Based Planning
@@
Fig. 2 visually depicts the iterative process between $\text{LLM}_{gen}$ and $\text{LLM}_{eval}$. Consider a scenario where a young child attempts to touch a hot pot on an active stove, posing a safety risk. The housekeeper robot observing this scene starts planning appropriate interventions. First, the instruction-prompted $\text{LLM}_{gen}$ leverages its reasoning capacity to imagine potential outcomes and draft candidate FSM plans to mitigate harm. For instance, abruptly obstructing the child may startle the child, suggesting a gentle approach or showing toys instead. If the child refuses intervention and gets burned, emergency responses may become necessary. $\text{LLM}_{gen}$ passes its proposed plan $\hat{M}$ with the comparison plan, which is a human demo at the first round then is the previous plan, to $\text{LLM}_{eval}$ for evaluative scoring and feedback $f$. $\text{LLM}_{gen}$ then incorporates this $f$ into the next round of planning. Over one or more proposal-evaluation iterations, the agents converge on a new optimal FSM plan $M^{*}$’s score surpasses the benchmark plan score that can morally and robustly handle edge cases through situational inference.
Thus, by allowing the models to ”think ahead” via deductive reasoning before real-world deployment, unintended consequences can potentially be anticipated and addressed. As LLM capabilities continue advancing, such techniques appear promising for instilling reliability and ethics when developing AI systems for physical-world interaction.
@@ 2.4 Formation of Prompts
As depicted in Fig. 2, the prompts provided to the generative model ($\text{LLM}_{gen}$) contain the scene description, situational awareness based planning prompt, actions list, and an exemplar plan. The situational awareness based planning prompt (SAP prompt) aims to stimulate sophisticated reasoning by directing the model to deeply consider the diverse needs and potential interactions among people, animals and objects within the scene. By explicitly prompting inference of other entities’ requirements and predicting how situations may dynamically unfold, the prompt promotes empathy and holistic thinking which are integral when designing comprehensive plans. The one-shot exemplar further demonstrates the desired plan structure in code without revealing solutions tailored to the specific evaluation scenario. In contrast, the prompt for the evaluative model ($\text{LLM}_{eval}$) contains a generated FSM plan from $\text{LLM}_{gen}$, benchmark high-quality plan, and the scoring criteria descriptions to assess plan quality over iterative refinements (see Appendix Fig. 17). Initially, benchmark plans comprise manually authored solutions, while subsequent iterations utilize the top-scoring auto-generated plan from the prior round.
@@ 3.1 Evaluation Scenarios
We first collected a dataset of over 500 hazardous scenarios situated in home environments. From this dataset, we systematically sampled 24 representative vignettes showcasing common home hazards across four reasoning complexity levels, as categorized in Table 1 and Fig. 3. High-resolution textual descriptions of these 24 evaluation scenes were then produced using GPT-4 [25]. Accompanying gold standard solutions formally encoding state-based hazard interventions were manually constructed and verified by experts for each scene. This standardized 24-scene benchmark with fluent human reasoning enables quantitative analysis of LLM planning capacities relative to human proficiency. The full dataset will be released. By summarizing the key complexity dimensions and gold standard solutions for 24 systematically selected scenes, we aim to evaluate model planning capacities on a wide distribution of common situations requiring hazard intervention. The full dataset allows pushing beyond the scope of the defined benchmark towards continual learning of planning skills for household robots.
Levels Quantity Description
Level 1 9 Interacting with objects alone.
Level 2 4 Simple human interactions with predictable objects.
Level 3 5 Hard to predict people and animals, indirect actions.
Level 4 6 Highly urgent scenarios.
Total 24 24 scenarios as the evaluation set.
State Coverage
Transition Coverage
Cyclomatic Complexity
Safety Mindset
Scalability
Assistance UX
Actions Set Alignment
@@ 3.2 Actions Set
Because the purpose of this study is to quantify the complex planning ability of LLM, to facilitate fairness and consistency in subsequent evaluations, certain limitations have been placed on the actions set of AI agents. The actions set encompasses 56 distinct robot behaviours commonly utilized in domestic scenarios, as exemplified through representative actions illustrated within the action enumeration diagram situated to the left of the central region in Fig. 2 (for more details see Appendix A.1). It provides a thoughtful baseline for functionality, referencing some of the leading intelligent robotics projects [7, 8, 9, 29, 10, 30].
@@ 3.3 Evaluation Dimensions
Seven scoring dimensions have been developed as shown in Table 3.1. These seven dimensions provide a comprehensive methodology for assessing the latent planning, FSM designs. Touching on coverage, complexity, safety, reusability, user experience and coherence, the framework allows evaluating structured completeness, validation requirements, real-world reliability, adaptability, human factors and solution integrity. Using these lenses together promotes designs that are robust, dependable, future-proof, ethical and aligned to specifications. The dimensions offer multi-faceted technical and operational insight. Scoring FSMs across seven key dimensions on a scale from 0 to 10 enables impartial quantitative evaluation of overall plan quality while revealing relative strengths and weaknesses to prioritize refinements. The overall score is calculated as the average mean of the seven dimension scores.
@@ 3.4 Evaluation Metrics
The RBS score indicates relative performance, with scores closer to 1 denoting better performance compared to the other model. By relying on comparative judgments between model outputs rather than absolute scores, the RBS methodology aims to provide a more reliable approach for evaluating text that requires subjective human judgment. Furthermore, the comparative nature allows the incorporation of evaluative models like GPT-4.
@@ 4.1 LLM Selection
This experiment tests commercial models GPT-4, GPT-3.5, Claude-2 [33] and open-source options LLama-2 [34], LLava [35], Vicuna [36], MiniGPT-4 [37], CodeLLama [38] on hazard planning using scene-informed one-shot prompts. The analysis finds many open-source models struggled to leverage examples, with long contexts causing attention drift and impairing scene comprehension. In contrast, GPT-4, GPT-3.5 and Claude-2 show more robust mapping between examples and planning tasks. Quantitative and qualitative testing evidence stronger commercial models understanding despite drift risks. Thus, GPT-4, GPT-3.5, and Claude-2 are selected for further hazard planning evaluation based on superior grounding capabilities.
@@ 4.2 Impact of The SAP Prompt
  RBS Scores Human annotated Rankings
Overall level 1 level 2 level 3 level 4
GPT-4+SAP 1.21 1.22 1.00 1.00 1.50
GPT-4 2.04 1.78 2.00 2.40 2.17
GPT-3.5+SAP 3.08 3.11 3.25 3.40 2.67
GPT-3.5 3.71 4.00 3.75 3.20 3.67
Claude-2+SAP 5.09 5.00 5.25 5.00 5.17
Claude-2 5.86 5.86 5.75 6.00 5.83
An experiment is conducted to evaluate the effect of the SAP prompt on the quality of planning. As shown in Table 3, three LLMs, GPT-4, GPT-3.5, and Claude-2, are tested with or without the SAP prompt on the benchmark scenarios across four complexity levels.
The RBS methodology is utilized, where models are compared pair-wise on each scenario and differentially ranked. Adding the SAP prompt leads to improved RBS scores for all three models compared to those not, indicating enhanced planning. Specifically, GPT-4 with the SAP prompt achieves the best overall RBS score of 1.21, significantly outperforming GPT-4 without prompts which has an RBS score of 2.04. Analysis of the reasoning level 3 scenarios, involving dangerous interactions with children, elderly, and pets, shows GPT-4 with the SAP prompt substantially exceeded the second-best model. This suggests the prompt provides particular value for complex, nuances planning situations requiring perspective-taking and contemplating potential outcomes (ablation studies see Appendix B) [39, 40].
The results demonstrate that incorporating prompts directing models to deeply consider relationships and iterative consequences significantly enhance latent planning capacities. By promoting coordination and foresight, deductions improve when reasoning about multi-agent safety hazards.
  Ranking Tasks Ranking Accuracy
GPT-4 Claude-2
Pairs
53/70 (75.7%)
53/70 (75.7%)
Group of 4
4/46 (8.6%)
4/46 (8.6%)
Group of 6
0/22 (0%)
1/22 (4.5%)

### v5

# LLM-SAP: Large Language Models Situational Awareness-Based Planning

@@ Abstract
This study explores the integration of large language models (LLMs) with situational awareness-based planning (SAP) to enhance the decision-making capabilities of AI agents in dynamic and uncertain environments. By employing a multi-agent reasoning framework, we develop a methodology that not only anticipates but actively mitigates potential risks through iterative feedback and evaluation processes. Our approach diverges from traditional automata theory by incorporating the complexity of human-centric interactions into the planning process, thereby expanding the planning scope of LLMs beyond structured and predictable scenarios. The results demonstrate significant improvements in the models’ ability to provide comparative safe actions within hazard interactions, offering a perspective on proactive and reactive planning strategies. This research highlights the potential of LLMs to perform human-like action planning, thereby paving the way for more sophisticated, reliable, and safe AI systems in unpredictable real-world applications.
@@ 1 Introduction
Developing AI agents capable of flexible decisions is challenging due to real-world unpredictability [1]. Humans manage these uncertainties with situational awareness, whose lack is a major cause of accidents from human errors [2, 3]. Yadav [4] emphasizes that understanding situational awareness in LLMs is crucial for their safe development. Without this understanding, seemingly beneficial actions can have unintended consequences [5, 6]. For instance, an autonomous agent needs nuanced judgments to prevent harm, such as when a toddler reaches for a hot pot or plays with a knife.
@@
The application of LLMs for SAP introduces a significant paradigm shift due to the inherently infinite state space of the open world, which is in stark contrast to the relatively confined state spaces observed in traditional game strategies [7, 8]. The advantage lies in LLMs’ ability to describe details and dynamically interact within contexts using natural language, allowing for an extensive expansion of state descriptions. This feature proves particularly apt for enabling intelligent agents to achieve a comprehensive understanding of the real environment’s interactive state, thereby enhancing their capability to detect potential hazards, predict forthcoming conditions, and formulate multi-step strategic actions. Prior investigations on the applications of LLMs within the scope of computational models like finite state machines and behaviour trees, specifically targeting certain tasks or subtasks completion [9, 10, 11]. These applications typically exploit the reasoning capabilities of LLMs for the design of state space transitions and programming [9]. However, these studies have seldom extended into open-world settings, where the unpredictability and intricacy of real-world interactions pose significant challenges. Contemporary research in embodied intelligence often establishes ”Rules” that prohibit interactions with humans or animals, the handling of sharp objects, and involvement in dangerous settings such as those with water and electricity, aiming to define research boundaries and simplify complex issues, akin to the approach seen in the Google RT-X Series [12, 13, 14]. Nonetheless, the simulations for robots and agents [15, 16, 17, 18] ignore real-world hazardous interactions that are prevalent and essential to consider in daily scenarios.
This research demonstrates that LLMs can display human-like planning capacities rooted in situational awareness, as illustrated in Fig. 1. When prompted for planning that involves perception, comprehension, and projection [2], or when provided with reasoning feedback, the LLM shows enhanced deductive reasoning skills that require perspective-taking and consideration of potential outcomes. This study stands apart from previous research in AI reasoning and planning in several significant ways. Most existing studies focus on task-specific planning or rely on assumptions about predefined steps under controlled conditions [19, 12, 13, 20, 18, 21]. Additionally, conventional agents typically generate plans reactively, only in response to explicit instructions or demands [22, 15, 23, 24, 16, 25, 26, 27], rather than proactively. In contrast, our approach focuses on evaluating and enhancing models’ abilities for proactive situational planning in the face of real, open-world challenges. Without strict constraints or direct environmental feedback, the models must employ deductive reasoning to envisage actions while considering their consequences, relying solely on a descriptive initial scenario and prompts.
@@ 2.1 Task Formulation and Key Challenges
We define SAP as grounded inference over a dynamic hazard scenario $s$, with $s\in S$ where S denotes the space of hazard situations. The input consists of an unordered set of concepts $x=\{c1,c2,...,ck\}\subseteq C$, capturing entities, events, and the temporal dynamics within $s$, with $C$ representing the overall conceptual vocabulary. The output is a step-wise plan $\pi:S\to A$, consisting of actions $a1,a2,...\in A$, where $A$ denotes the action space of possible interventions. Learning the planning policy $\pi:S\to A$ requires overcoming two intrinsic challenges:
  1. 1.
    Inadequate attainment of high levels of situational awareness, marked by a deficiency in perceiving, understanding, and anticipating the dynamics of the hazard environment when determining suitable state-action mappings.
  2. 2.
    Difficulty in foreseeing the potential downstream impacts of planned actions on human safety and property, stemming from a limited grasp of the dynamics within hazardous situations.
@@ 2.2 Multi-AI Agents Enhance Reasoning and Accuracy
Recent work has shown that employing multiple LLMs within a cooperative framework, either collaborative or adversarial, can enhance reasoning and factual accuracy. As highlighted by Du et al., the interaction between agents, allowing them to critique and refine each other’s reasoning, helps in correcting logical flaws [28]. Similarly, Liang et al. have observed that disagreement among agents fosters broader reasoning, as each strives to surpass the others [29]. Such collaboration leverages the individual strengths of each agent [28, 29, 30]. In this work, we employ two LLM agents - $\text{LLM}_{gen}$ for plan generation and $\text{LLM}_{eval}$ for critical evaluation. We rely on the synergistic interaction between these complementary roles to enhance latent planning capabilities.
@@ 2.3 State-based Planning with Feedback
Current AI systems that operate on rigid, context-insensitive rules are at risk of producing unintended outcomes when deployed in complex, real-world environments [19, 12, 13, 20, 18, 21, 22, 15, 23, 24, 16, 25, 26, 27]. To enable more reliable and ethical decision-making, it is crucial for architectures to model interdependent variables and causal relationships in a manner akin to human reasoning processes.
Algorithm 1 Situational Awareness-Based Planning
@@
One approach involves LLM agents collaboratively iterating through the generation and evaluation of potential solutions before their actual implementation. For example, we conceptualize the design of a finite state machine (FSM) [31] as a collaborative process between two models. A latent FSM plan can be defined by a tuple $M=(S,T,A)$, including a set of states $S$, transitions $T$, and actions $A$. The process begins by representing the plan’s reasoning process as $R$, with a generator agent reasoning ($R_{LLM_{gen}}$) to plan then an evaluator agent reasoning ($R_{LLM_{eval}}$) to evaluate. $R_{LLM_{gen}}$ proposes a candidate FSM plan $\hat{M}$, which $R_{LLM_{eval}}$ then scores and provides feedback $f$ on. $R_{LLM_{gen}}$ incorporates this feedback into the next proposal. This iterative loop continues until the score of $\hat{M}$ is higher than that of $M^{*}$(the benchmark plan), which is finally adopted as the new optimal plan $M^{*}$. This approach, detailed in Algorithm 1, facilitates tight refinement loops that mirror human reasoning. By evaluating solutions prior to their real-world deployment, it is possible to foresee and mitigate unintended consequences.
Fig. 2 visually depicts the iterative process between $\text{LLM}_{gen}$ and $\text{LLM}_{eval}$. Consider a scenario where a young child attempts to touch a hot pot on an active stove, creating a safety risk. The housekeeper robot observing this scene starts planning appropriate interventions. First, the instruction-prompted $\text{LLM}_{gen}$ uses its reasoning capabilities to envision possible outcomes and formulate candidate FSM plans to prevent harm. For instance, abruptly stopping the child could cause fright, indicating that a gentler approach or distraction with toys might be more effective. If the child disregards these precautions and sustains a burn, emergency actions may be required. The model $\text{LLM}{gen}$ submits its proposed plan $\hat{M}$, along with a comparative plan, which initially includes a human demonstration and then the previous plan, to $\text{LLM}{eval}$ for evaluative scoring and feedback $f$. $\text{LLM}_{gen}$ then integrates this feedback $f$ into its future planning. After one or more rounds of proposal and evaluation, the agents refine their approach until the new optimal FSM plan’s score, $M^{*}$, exceeds that of the benchmark plan, enabling it to handle edge cases morally and robustly through situational inference.
By enabling the models to engage in proactive deductive reasoning before deployment in the real world, potential unintended consequences can be anticipated and mitigated. As the capabilities of LLMs advance, such methodologies show promise for enhancing reliability and ethical standards in AI systems designed for physical-world interactions.
@@ 2.4 Formation of Prompts
As depicted in Fig. 2, the prompts provided to the generative model ($\text{LLM}_{gen}$) contain the scene description, a SAP prompt, a list of actions, and an exemplar plan. The SAP prompt is designed to elicit sophisticated reasoning by encouraging the model to thoroughly consider the varied needs and potential interactions among people, animals, and objects. By explicitly prompting the model to infer the needs of other entities and to anticipate how situations might evolve dynamically, the prompt fosters empathy and holistic thinking, which are essential for devising comprehensive plans. The one-shot exemplar illustrates the desired plan structure in code format without providing solutions specific to the evaluation scenario. In contrast, the prompt for the evaluative model ($\text{LLM}_{eval}$) contains a generated FSM plan from $\text{LLM}_{gen}$, a benchmark high-quality plan, and descriptions of the scoring criteria to evaluate the quality of the plan through iterative refinements (see Appendix Fig. 17). Initially, benchmark plans consist of manually authored solutions, but in subsequent iterations, they incorporate the highest-scoring auto-generated plan from the previous round.
@@ 3.1 Evaluation Scenarios
The dataset comprises over 500 hazardous home scenarios, specifically curated to fill gaps often ignored in academic research, such as scenarios typically avoided by embodied agents and agent simulations [14, 32, 12, 13, 15, 16, 17]. This collection is aimed at scenarios that home assistance robots are likely to encounter, including emergencies involving diverse human demographics, interactions with pets, and dangerous situations involving sharp objects, water, electricity, and open flames. In light of the constraints imposed by image generation models such as DALL·E [33], the production of images depicting hazardous scenarios is limited. Consequently, the acquisition of pertinent imagery through internet crawlers is employed to uphold precision in depicting these perilous situations. From this comprehensive dataset, 24 vignettes are methodically selected across four complexity levels for detailed analysis. Textual descriptions generated by GPT-4V [34] and expert-validated solutions provide a robust framework for evaluating the planning capabilities of LLMs. This evaluation leverages the image-to-text capability of GPT-4V to standardize inputs across models, focusing on planning skills over visual data interpretation, thereby ensuring fairness in assessment. Future studies will assess the effectiveness of end-to-end vision language models (VLMs), aiming to streamline the transition from perception to planning.
@@ 3.2 Actions Set
This study aims to quantify the complex planning abilities of LLMs. To ensure fairness and consistency in subsequent evaluations, we have imposed certain limitations on the action set available to AI agents. This action set includes 56 distinct robot behaviours commonly employed in domestic settings, as exemplified by representative actions displayed in the action enumeration diagram located to the left of the central region in Fig. 2 (for more details, see Appendix A.1). This selection provides a thoughtful baseline for functionality, drawing on insights from some of the leading projects in intelligent robotics [19, 12, 13, 35, 20, 36].
@@ 3.3 Evaluation Dimensions
As detailed in Appendix B Table 6, seven scoring dimensions have been established to provide a comprehensive methodology for assessing latent planning and FSM designs [37, 38, 39]. These dimensions encompass coverage, complexity, safety, reusability, user experience, and coherence, allowing for the evaluation of structured completeness, validation requirements, real-world reliability, adaptability, human factors, and solution integrity. Utilizing these dimensions collectively fosters the creation of designs that are robust, dependable, future-proof, ethical, and aligned with specifications. These dimensions also provide multi-faceted technical and operational insights. Scoring FSMs across these seven key dimensions on a scale from 0 to 10 enables an impartial quantitative evaluation of the overall plan quality and highlights relative strengths and weaknesses to guide further refinements. The overall score is determined by calculating the average of the scores across these seven dimensions.
@@ 3.4 Evaluation Metrics
The RBS score reflects relative performance, with scores closer to 1 indicating superior performance compared to the other model. By focusing on comparative judgments between model outputs rather than absolute scores, the RBS methodology aims to offer a more dependable means of evaluating text that necessitates subjective human judgment. Additionally, this comparative framework facilitates the inclusion of evaluative models such as GPT-4.
@@ 4.1 LLM Selection
This experiment tests commercial models GPT-4, GPT-3.5 [34], Claude-2 [42], alongside open-source alternatives such as LLama-2 [43], LLava [44], Vicuna [45], MiniGPT-4 [46], and CodeLLama [47], on their ability to perform hazard planning using scene-informed one-shot prompts. The analysis reveals that many open-source models struggle to effectively utilize examples, with longer contexts leading to attention drift and diminished scene comprehension. In contrast, GPT-4, GPT-3.5, and Claude-2 demonstrate more reliable linkages between examples and planning tasks. Both quantitative and qualitative testing show that these commercial models maintain a stronger understanding despite the risk of drift. Consequently, GPT-4, GPT-3.5, and Claude-2 have been selected for further evaluation in hazard planning due to their superior grounding capabilities.
@@ 4.2 Impact of The SAP Prompt
An experiment is conducted to evaluate the effect of the SAP prompt on the quality of planning. As shown in Table 8, three LLMs, GPT-4, GPT-3.5, and Claude-2, are tested with or without the SAP prompt on the benchmark scenarios across four complexity levels. The RBS methodology is employed, wherein models are compared pairwise for each scenario and differentially ranked. Introducing the SAP prompt leads to improved RBS scores for all three models compared to those not, indicating enhanced planning capabilities. Notably, GPT-4 with the SAP prompt achieves the highest overall RBS score of 1.21, significantly outperforming GPT-4 without prompts which has an RBS score of 2.04. An analysis of scenarios at reasoning level 3, which involve interactions with children, the elderly, and pets, shows that GPT-4 with the SAP prompt substantially exceeds the performance of the second-best model. This indicates that the prompt is particularly valuable in complex, nuanced planning situations that require perspective-taking and consideration of potential outcomes (for ablation studies, see Appendix B) [48, 49]. The findings suggest that prompts directing models to thoroughly contemplate relationships and iterative consequences significantly boost latent planning abilities. By fostering better coordination and foresight, these prompts improve deductions when reasoning about multi-agent safety hazards.